# Two axes in non-associative algebras with a Frobenius form

### v5

@@ 1. Introduction
Throughout this paper $A$ is a commutative non-associative algebra over a field $\mathbb{F}$ of characteristic not $2.$ Assume that (,)(\ ,\ ) is a Frobenius form on $A,$ that is, (,)(\ ,\ ) is a symmetric bilinear map such that $(xy,z)=(x,yz),$ $\forall x,y,z\in A.$ We denote by $\mathfrak{R}$ the radical of (,)(\ ,\ ). Recall that for $z\in A,\ \lambda\in\mathbb{F},$ we denote $A_{\lambda}(z)=\{x\in A\mid zx=\lambda x\}.$ We direct the reader to the definition 1.1 below and the beginning of §2 for the definitions required for the introduction.
@@
    $\Big(b_{\frac{1}{2}}(y_{\frac{1}{2}}x_{0})\Big)_{0}=(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{0}-\Big((b_{\frac{1}{2}}x_{0})y_{\frac{1}{2}}\Big)_{0}.$
@@ 2. The polynomials $Q_{c}(x)$ and $P_{c}(x,y)$
$(b)\implies(a):$ Assume that $(b)$ holds. Since $(c,c)=1,$ $Q_{c}(c)=0.$ Let $x_{0}\in A_{0}(c)$ and $x_{\frac{1}{2}}\in A_{\frac{1}{2}}(c).$ By Lemma 2.1, $(c,x_{0})=(c,x_{\frac{1}{2}})=0.$ Now $Q_{c}(x_{0})=c(cx_{0})-\frac{1}{2}(cx_{0})-\frac{1}{2}(c,x_{0})c)=0,$ and $Q_{c}(x_{\frac{1}{2}})=c(cx_{\frac{1}{2}})-\frac{1}{2}(cx_{\frac{1}{2}}+(c,x_{\frac{1}{2}})c)=\frac{1}{4}x_{\frac{1}{2}}-\frac{1}{4}x_{\frac{1}{2}}-\frac{1}{2}(c,x_{\frac{1}{2}})c=0.$
@@ 3. Derivations
(2) By Lemma 3.2(2)
$\textstyle{\Big(\big([L_{a},L_{v}](x)\big)y_{\frac{1}{2}}\Big)_{\frac{1}{2}}=(\frac{1}{2}(v,x)a-\frac{1}{2}vx_{\frac{1}{2}})y_{\frac{1}{2}}}$
$=\textstyle{\frac{1}{4}(v,x)y_{\frac{1}{2}}-\frac{1}{2}(vx_{\frac{1}{2}})y_{\frac{1}{2}}.}$
@@ 4. Obtaining identities in AA
$\textstyle{=w+\frac{1}{2}(a,b)(1-2(a,b))x_{\frac{1}{2}}+(b,x_{\frac{1}{2}})b_{\frac{1}{2}}+(1-2(a,b))x_{\frac{1}{2}}b_{0}}$
@@
$\textstyle{=w+\frac{1}{2}(a,b)(1-2(a,b))x_{\frac{1}{2}}+\frac{1}{2}(b,x_{\frac{1}{2}})b_{\frac{1}{2}}+(1-2(a,b))x_{\frac{1}{2}}b_{0}}$
@@
$+(b,x_{\frac{1}{2}})\Big(b_{\frac{1}{2}}^{2}\Big)_{0}-2\Big((b_{0}x_{\frac{1}{2}})b_{\frac{1}{2}}\Big)_{0}.$
@@
    $b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})+(a,b)(b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+2((x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})$
@@
$=(x_{\frac{1}{2}}b_{0})y_{0}+(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}-2\Big(\frac{1}{2}(a,b)x_{\frac{1}{2}}(b_{\frac{1}{2}}y_{0})\Big)$
@@
$=(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}-(a,b)x_{\frac{1}{2}}(b_{\frac{1}{2}}y_{0})-2(x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})-2(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0}).$
@@
$\textstyle{-2\Big(\frac{1}{2}(a,b)y_{\frac{1}{2}}(b_{\frac{1}{2}}x_{0})+(b_{\frac{1}{2}}x_{0})(y_{\frac{1}{2}}b_{0})+(b_{\frac{1}{2}}x_{0})(b_{\frac{1}{2}}y_{\frac{1}{2}})\Big).}$
@@
(3) This follows from comparing the 0 components of both sides of (2). ∎
@@
    Interchanging $x_{\frac{1}{2}}$ and $y_{\frac{1}{2}}$ in (2) we get
@@
      $(1-(a,b))(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}+2b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))+2((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}$
@@
$b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})-(a,b)b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})-2b_{\frac{1}{2}}((y_{\frac{1}{2}}b_{0})x_{\frac{1}{2}})-(a,b)(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}-2((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}$
$=\quad(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}-(a,b)x_{\frac{1}{2}}(b_{\frac{1}{2}}y_{\frac{1}{2}})$
$-2(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})-(a,b)(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}-2(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{\frac{1}{2}}b_{0})$
$\iff$
$b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})-(a,b)b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})-2b_{\frac{1}{2}}((y_{\frac{1}{2}}b_{0})x_{\frac{1}{2}})-2((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}$
$=\quad(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}-2(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})-(a,b)(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}-2(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{\frac{1}{2}}b_{0}).$
@@
$\iff\Big(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})\Big)_{0}+2(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))b_{0}+2\Big(b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})\Big)_{0}$
@@
    $(a,x_{\frac{1}{2}}b_{0},y_{\frac{1}{2}})=\frac{1}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}}).$
@@
(1) By Theorem 4.9(1),
$(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}}))b_{0}=\Big(\Big[(a,x_{\frac{1}{2}}y_{\frac{1}{2}})a+\Big(x_{\frac{1}{2}}y_{\frac{1}{2}}\Big)_{0}\Big]b_{\frac{1}{2}}\Big)b_{0}$
$\textstyle{=\frac{1}{4}(1-(a,b))(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+\frac{1}{4}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.}$
(2) Replacing $y_{\frac{1}{2}}$ with $y_{\frac{1}{2}}b_{0}$ in (1), and using Theorem 4.9(1), we see that
$\textstyle{(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))b_{0}=\frac{1}{4}(1-(a,b))(a,x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))b_{\frac{1}{2}}+\frac{1}{4}(b_{0},x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))b_{\frac{1}{2}}}$
$=\textstyle{\frac{1}{4}(1-(a,b))(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.}$
(3) is clear and (4) follows from Theorem 4.9(1). (5) follows from (3) and (4) replacing $y_{\frac{1}{2}}$ with $y_{\frac{1}{2}}b_{0}.$ Similarly (6) follows from (4). For (7) we have $(b,(x_{\frac{1}{2}}b_{0})y_{\frac{1}{2}})=\frac{1}{2}(a,b)(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})+\frac{1}{2}(1-(a,b))(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})=\frac{1}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}}),$ and (8) follow from (7).
@@
  9. (9)
    If $(a,b)\notin\{0,1,\frac{1}{4}\},$ then,
    $b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})+(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}\Big]\\ =\frac{1}{2}(b_{\frac{1}{2}},y_{\frac{1}{2}})x_{\frac{1}{2}}+\frac{1}{2}(b,x_{\frac{1}{2}})y_{\frac{1}{2}}+(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.$
@@
$-\frac{1}{2}(a,b)(b,y_{\frac{1}{2}})x_{\frac{1}{2}}-(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}-(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{\frac{1}{2}}-\frac{1}{2}(a,b)(b,x_{\frac{1}{2}})y_{\frac{1}{2}}-(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}-(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{\frac{1}{2}}$
@@
$+2(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0})))b_{0}+2(b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})))b_{0}+(1-(a,b))\Big[((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}+2((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}\Big]$
@@
$\Big[((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}+2((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}\Big]$
@@
$(1-(a,b))(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})$
$+\Big[((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}+2((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}\Big]$
@@
$\textstyle{\frac{1}{2}\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}\Big]-\Big[((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}+2((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}\Big]}$
$\textstyle{-\frac{1}{4}(1-(a,b))(a,x_{\frac{1}{2}}y_{\frac{1}{2}})+\frac{1}{4}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}}).}$
(5) Putting $x_{\frac{1}{2}}b_{0}$ in place of $x_{\frac{1}{2}}$ in equation (4.1) we get
$\Big[(b,y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})b_{0}+(b,(x_{\frac{1}{2}}b_{0}))y_{\frac{1}{2}}b_{0}\Big]-2\Big[((b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0}))b_{0}+((b_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))y_{\frac{1}{2}})b_{0}\Big]$
@@
(7) (a) Replacing in (6), $x_{\frac{1}{2}},y_{\frac{1}{2}}$ with $x_{\frac{1}{2}}b_{0},y_{\frac{1}{2}}b_{0},$ using Lemma 4.12((8)&(10)), we get
@@
$(2(a,b)-1)(1-(a,b))\Big((x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0})\Big)_{0}+(2(a,b)-1)(1-(a,b))\Big((y_{\frac{1}{2}}b_{0})(x_{\frac{1}{2}}b_{0})\Big)_{0}$
@@
(8) Replacing in (6) $x_{\frac{1}{2}}$ with $x_{\frac{1}{2}}b_{0},$ using Lemma 4.9((9)&(7)) and (5) we get
@@
(9) Since $(a,b)\notin\{1,\frac{1}{4}\},$ we can use Theorem 7.1 to get
$2b_{\frac{1}{2}}\Big[(\frac{1}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})a+\Big((x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0})\Big)_{0}\Big]+2b_{\frac{1}{2}}\Big[(\frac{1}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})a+\Big((y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})\Big)_{0}\Big]$
$(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+2b_{\frac{1}{2}}\Big[\Big(y_{\frac{1}{2}}b_{0})x_{\frac{1}{2}}\Big)_{0}+\Big((x_{\frac{1}{2}}b_{0})y_{\frac{1}{2}}\Big)_{0}\Big]$
$(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+2b_{\frac{1}{2}}(\Big(x_{\frac{1}{2}}y_{\frac{1}{2}}\Big)_{0}b_{0})$
$(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+(1-(a,b))b_{\frac{1}{2}}(\Big(x_{\frac{1}{2}}y_{\frac{1}{2}}\Big)_{0})+\frac{1}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}$
$\textstyle{\frac{3}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})+(1-(a,b))b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})-\frac{1}{2}(1-(a,b))(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.}$
Hence we get from (6),
$\textstyle{+\frac{1}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}-\frac{1}{2}(1-(a,b))(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+(b,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}}$
$\textstyle{\frac{3}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})+(1-(a,b))b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})-\frac{1}{2}(1-(a,b))(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}}$
$\iff$
$-(a,b)b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})-(a,b)\Big[(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}\Big]$
$\textstyle{+\frac{1}{2}(a,b)\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}\Big]+(a,b)(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}=0.}$

### v6

@@ 1. Introduction
Throughout this paper $A$ is a commutative non-associative algebra over a field $\mathbb{F}$ of characteristic not $2.$ Assume that (,)(\ ,\ ) is a Frobenius form on $A,$ that is, (,)(\ ,\ ) is a symmetric bilinear map such that $(xy,z)=(x,yz),$ $\forall x,y,z\in A.$ We denote by $\mathfrak{R}$ the radical of (,)(\ ,\ ). Recall that for $z\in A,\ \lambda\in\mathbb{F},$ we denote $A_{\lambda}(z)=\{x\in A\mid zx=\lambda x\}.$ We direct the reader to Definition 1.1 below and the beginning of §2 for the definitions required for the introduction.
@@
    $\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}=(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}-\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}.$
@@ 2. The polynomials $Q_{c}(x)$ and $P_{c}(x,y)$
$(b)\implies(a):$ Assume that $(b)$ holds. Since $(c,c)=1,$ $Q_{c}(c)=0.$ Let $x_{0}\in A_{0}(c)$ and $x_{\frac{1}{2}}\in A_{\frac{1}{2}}(c).$ By Lemma 2.1, $(c,x_{0})=(c,x_{\frac{1}{2}})=0.$ Now $Q_{c}(x_{0})=c(cx_{0})-\frac{1}{2}(cx_{0})-\frac{1}{2}(c,x_{0})c)=0,$ and $Q_{c}(x_{\frac{1}{2}})=c(cx_{\frac{1}{2}})-\frac{1}{2}cx_{\frac{1}{2}}+(c,x_{\frac{1}{2}})c)=\frac{1}{4}x_{\frac{1}{2}}-\frac{1}{4}x_{\frac{1}{2}}-\frac{1}{2}(c,x_{\frac{1}{2}})c=0.$
@@ 3. Derivations
(2) By Lemma 3.2(2),
$\Big(\big([L_{a},L_{v}](x)\big)y_{\frac{1}{2}}\Big)_{\frac{1}{2}}=(\frac{1}{2}(v,x)a-\frac{1}{2}vx_{\frac{1}{2}})y_{\frac{1}{2}}=\frac{1}{4}(v,x)y_{\frac{1}{2}}-\frac{1}{2}(vx_{\frac{1}{2}})y_{\frac{1}{2}}.$
@@ 4. Obtaining identities in AA
$\textstyle{=w+\frac{1}{2}(a,b)(1-(a,b))x_{\frac{1}{2}}+(b,x_{\frac{1}{2}})b_{\frac{1}{2}}+(1-2(a,b))x_{\frac{1}{2}}b_{0}}$
@@
$\textstyle{=w+\frac{1}{2}(a,b)(1-(a,b))x_{\frac{1}{2}}+\frac{1}{2}(b,x_{\frac{1}{2}})b_{\frac{1}{2}}+(1-2(a,b))x_{\frac{1}{2}}b_{0}}$
@@
$+(b,x_{\frac{1}{2}})\Big(b_{\frac{1}{2}}^{2}\Big)_{0}-2\Big((b_{0}x_{\frac{1}{2}})b_{\frac{1}{2}}\Big)_{0},$
@@
    $b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})+(a,b)(b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+2(x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})$
@@
$=\textstyle{(x_{\frac{1}{2}}b_{0})y_{0}+(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}-2\Big(\frac{1}{2}(a,b)x_{\frac{1}{2}}(b_{\frac{1}{2}}y_{0})\Big)}$
@@
$=(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}-(a,b)x_{\frac{1}{2}}(b_{\frac{1}{2}}y_{0})-2(x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})-2(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0}).\qed$
@@
$\textstyle{-2\Big(\frac{1}{2}(a,b)y_{\frac{1}{2}}(b_{\frac{1}{2}}x_{0})+(b_{\frac{1}{2}}x_{0})(y_{\frac{1}{2}}b_{0})+(b_{\frac{1}{2}}x_{0})(b_{\frac{1}{2}}y_{\frac{1}{2}})\Big)}$
@@
(3) This follows from comparing the $1$ plus 0 components of both sides of (2). ∎
@@
    Interchanging $x_{\frac{1}{2}}$ and $y_{\frac{1}{2}}$ in (2) we get:
@@
      $(1-(a,b))(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}+2b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))+2((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}.$
@@
$(1-(a,b))b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})-2b_{\frac{1}{2}}((y_{\frac{1}{2}}b_{0})x_{\frac{1}{2}})-2((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}$
$=\quad(1-(a,b))(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}-2(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})-2(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{\frac{1}{2}}b_{0}).$
@@
    $(a,(x_{\frac{1}{2}}b_{0})y_{\frac{1}{2}})=\frac{1}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}}).$
@@
(1) By Theorem 4.9(1), $(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}}))b_{0}=\Big(\Big[(a,x_{\frac{1}{2}}y_{\frac{1}{2}})a+\Big(x_{\frac{1}{2}}y_{\frac{1}{2}}\Big)_{0}\Big]b_{\frac{1}{2}}\Big)b_{0}\\ \textstyle{=\frac{1}{4}(1-(a,b))(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+\frac{1}{4}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.}$
(2) Replacing $y_{\frac{1}{2}}$ with $y_{\frac{1}{2}}b_{0}$ in (1), and using Theorem 4.9(1), we see that ${(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))b_{0}=\frac{1}{4}(1-(a,b))(a,x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))b_{\frac{1}{2}}+\frac{1}{4}(b_{0},x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))b_{\frac{1}{2}}}\\ =\textstyle{\frac{1}{4}(1-(a,b))(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.}$
(3) is clear and (4) follows from Theorem 4.9(1). (5) follows from (3) replacing $y_{\frac{1}{2}}$ with $y_{\frac{1}{2}}b_{0}.$ Similarly (6) follows from (4). For (7) we have $(b,(x_{\frac{1}{2}}b_{0})y_{\frac{1}{2}})=\frac{1}{2}(a,b)(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})+\frac{1}{2}(1-(a,b))(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})=\frac{1}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}}),$ and (8) follow from (7).
@@
    $2b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))+2b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})).$
@@
$+2(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0})))b_{0}+2(b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})))b_{0}+(1-(a,b))\Big[((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}+((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}\Big]$
@@
$\Big[((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}+((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}\Big]$
@@
$(1-(a,b))(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}$
$+\Big[((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}+((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}\Big]$
@@
$\textstyle{\frac{1}{2}\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}\Big]-\Big[((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}+((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}\Big]}$
$\textstyle{-\frac{1}{4}(1-(a,b))(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+\frac{1}{4}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.}$
(5) Putting $x_{\frac{1}{2}}b_{0}$ in place of $x_{\frac{1}{2}}$ in equation (4.1) we get (using also Lemma 4.12(9)),
$\Big[(b,y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})b_{0}+(b,x_{\frac{1}{2}}b_{0})y_{\frac{1}{2}}b_{0}\Big]-2\Big[((b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0}))b_{0}+((b_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))y_{\frac{1}{2}})b_{0}\Big]$
@@
(7) (a) Replacing in (6), $x_{\frac{1}{2}},y_{\frac{1}{2}}$ with $x_{\frac{1}{2}}b_{0},y_{\frac{1}{2}}b_{0},$ using Lemma 4.12 (5), (6), (8) and (10), we get
@@
$(2(a,b)-1)(1-(a,b))\Big((x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0})\Big)_{0}$
@@
(8) Replacing in (6) $x_{\frac{1}{2}}$ with $x_{\frac{1}{2}}b_{0},$ using Lemma 4.12 and part (7) we get